Week of 180618

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Kate (chair, WLCG), Julia (WLCG), Maarten (ALICE, WLCG), Andrea V. (LHCb), Miro P. (DB), Nicolo (ATLAS), Borja (monit), Roberto (storage), Andrei (DB)
  • remote: Andrew (NL-T1), Xavier (KIT), Marcelo (CNAF), Sang Un (KISTI), Christian (NDGF), Christoph (CMS), Di (TRIUMF), Dave M (FNAL), Xin (BNL), David (IN2P3)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Nicoḷ Magini reporting as ATLAS CRC for last week - no CRC this week
    • Stable operation with peaks > 350k slots including Sim@P1 resources during MD/TS
      • Current reprocessing campaign might cause heavy load on squid infrastructure, looks OK for now, keeping an eye on it.
      • Currently high job failure rate on INFN-T1_MCORE queue with input read error GGUS:135692 - possibly files on gpfs are not accessible from WNs ATM?
    • Transfers: 11 GB/s
      • Issue with high request load on FZK-LCG2 tape for staging of input for reprocessing GGUS:135651 - ongoing work by FZK and ATLAS people to make the queue more sustainable
      • INFN-T1 storage issue for few hours on 13/06 GGUS:135658 - fixed
      • On 15/6 BNL FTS was unresponsive for two hours - came back before ticket submitted.
      • A couple of issues with transfer submission on Rucio side (problem with urllib dependency, expired proxy delegation, slow daemon), fixed.
      • According to space monitoring FZK-LCG2_SCRATCHDISK is full of dark data after deletion on Fri 15/6 - either deletion left behind dark data, or space reporting is inconsistent.
    • Started tape carousel tests - first round done at BNL, now involving other T1s
Julia mentioned it is possible to compare accounting numbers with WLCG accounting portal for the FZK issue.

  • CMS reports ( raw view) -
    • CPU usage remains quite high, averaging ~280k cores (~220k for production) this week
    • The cleanup after the accidental deletion of data at T0 continues. T1 sites were asked to delete files on tape that were replaced by newly created ones after the cleanup. There were many GGUS tickets for this -- some are completed (GGUS:135652, GGUS:135655), while others are still open (GGUS:135653 (PIC), GGUS:135654 (IN2P3), GGUS:135656 (FNAL)).
    • GGUS:135673, problems with SAM tests at CNAF, comments on the ticket imply the trouble is past but the ticket is still open.

  • ALICE -
    • Normal to high activity on average
    • CERN: EOS-ALICE was down for 2.5h on Wed
    • CERN: alien.web.cern.ch down since Fri afternoon
      • update after the meeting: it was not due to the Drupal DB problem (OTG:0044511)
      • instead, the problem was cured by clearing the caches on the web server

  • LHCb reports ( raw view) -
    • Activity
      • Data reconstruction for 2018 data, MC simulation, user jobs
    • Site Issues

Sites / Services round table:

  • ASGC: nc
  • BNL: FTS was unresponsive for 2 hours, job submission was also blocked. A network issue outside BNL is suspected.
  • CNAF:
    • LHCb 2 ticket opened as mentioned,
      • transfer failures, we had an issue with SRM last week that Affected Atlas and LHCb, the frontend was restarted and seems normal now, it is under investigation to understand the cause.
      • Few jobs running, there was a major issue affecting all CEs and all VOs that was solved friday afternoon, after normalization the jobs built up again and we have now 10K LHCb jobs running
    • For Atlas, the issues is still under investigation, also probably related with the Storm frontend.
    • Last friday, CNAF was used as a attack amplifier during a moment when ACL was down for modification. 4 Machines were used for a mirror-attack. No sign of system compromising yet, it's under investigation the source of the attack and/or any possible consequence in our system
  • EGI: nc
  • FNAL: NTR
  • IN2P3: NTR
  • JINR: nc
  • KISTI: NTR
  • KIT: One border router needed a reboot on Thursday night for a yet unknown reason. Two line cards in it did not come back online after that. Hosts plugged into those were moved to other cards. The dCache core services for ATLAS and LHCb did not recover from this incident and needed to be restarted Friday morning.
  • NDGF: Upgraded site to dCache 3.2 this morning. Site is back in production.
  • NL-T1: Issues with data transfer. Increase of the no of SRM threads solved the issues.
  • NRC-KI: nc
  • OSG: nc
  • PIC: nc
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services: nc
  • CERN storage services: EOSATLAS update started today at 14:30 (will last ~1h)
  • CERN databases:
    • The latest OS updates are being applied these weeks to the hosts where the database services are running. This intervention is transparent and database services are available at all times.
  • GGUS: NTR
  • Monitoring:
    • Final reports for the May availability sent around
  • MW Officer: nc
  • Networks: nc
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2018-06-18 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback